The 8.50-cm-long second hand on a watch rotates smoothly.

What is the speed of the tip of the hand?

please include correct units
Answer #1

speed of the tip of the hand= tangential velocity = angular velocity * radius

First, we need to find teh angular velocity:

Angular velocity basically means the number of radians covered in a second

The second hand goes through 2\pi radians in 60secs

so, angular velocity = 2\pi/60 = 0.105rad/sec

Therefore, tangential velocity = 0.105*8.5*10-2 =8.9*10-3m/s